OpenStudy (anonymous):

how do you solve for y in both equations ? 1/x +1/y =3 and 1/x +2/y +3/y=4 in the fastest way possible ? :D

OpenStudy (anonymous):

u cud simply subtract both the equations!!

OpenStudy (anonymous):

And you get \[ \frac{3}{y}+\frac{2}{y}-\frac{1}{y}=4 -3 \]

OpenStudy (anonymous):

explain please! :D no i meant separately ....

OpenStudy (anonymous):

\[ \frac 4 y = 1\\ y=4 \]

OpenStudy (anonymous):

Replace in one of the equation and solve for x

OpenStudy (anonymous):

\[ \frac{1}{x}+\frac{1}{4}=3\\ x=\frac 4{11} \]

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yup, thank youu!

OpenStudy (anonymous):

yw